the 2017-2018 annual rainfall in Los Angeles was 4.79 inches (that is the water falling over the course of a year would cover the ground to a depth of 4.79 inches without drainage). if the area of my house is 26 acres what mass ( in metric tons, tonne) of water fell on the house in 2017-2108? (use the density of water at 25 degrees celsius= 0.997g/mL )

640acre=1mi^2

1tonne=1000kg
Answer #1

1 inch = 2.54 cm

1 ml = 1 cm3

1 acre = 4.047*10^7

4.79 inch = 12.16 cm

26 acre = 1.052*10^9cm2

volume of water = area*height

= 1.052*10^9*12.16

= 12.798*10^9

mass of water = volume * density

= 12.798*10^9*0.997 gms

= 12.76*10^9 gms

mass of water in tonne = 12.76*10^3 tonne

please like the solution